OpenStudy (akashdeepdeb):

PLEASE TELL ME THE CONCEPT: We can say that Instantaneous acceleration is The derivative of velocity w.r.t time = \(\frac{dv}{dt}\) But what does it mean when we say that instantaneous acceleration can also mean The second derivative of x w.r.t time = \(\frac{d^2x}{dt^2}\) ? I am new to calculus!! Please tell me what this means PRACTICALLY! :)

OpenStudy (rsadhvika):

velocity is the derivatve of x w.r.t time

OpenStudy (akashdeepdeb):

Okay.

OpenStudy (akashdeepdeb):

I know that but then what? :D

OpenStudy (rsadhvika):

\(v = \frac{dx}{dt}\) \(a= \frac{dv}{dt}\) combine them both

OpenStudy (akashdeepdeb):

Like this? \[a = \frac{ dv }{ dt } = \frac{ d \frac{ dx }{ dt } }{ dt }\] So we can just multiply the d and the dt like that directly?

OpenStudy (akashdeepdeb):

So would I get \[a = \frac{ d^2x }{ d^2t^2} ?\]

OpenStudy (rsadhvika):

\(a= \frac{dv}{dt} = \frac{d}{dt} (\frac{dx}{dt})\) think of d/dt as an operation. taking derivative of some quantity w.r.t. time

OpenStudy (akashdeepdeb):

But but what DOES IT MEAN?

OpenStudy (akashdeepdeb):

From what I understand v = dx/dt Means as time becomes really small dx = small change in distance dt = small change in time. WHat does, what you wrote, mean??

OpenStudy (rsadhvika):

okay you're mixing two different but related questions :- 1) wat does the *notation* \(\frac{d^2x}{dt^2}\) mean 2) wat does it mean wid regards to acceleration/velocity/distance

OpenStudy (akashdeepdeb):

I have no clue. :/ TEACH ME. :)

OpenStudy (rsadhvika):

answering first question is simple. that notation is, as the name implies is just a notation which refers to "taking derivative of x two times"

OpenStudy (akashdeepdeb):

And what does that mean? Taking the derviative twice? :/

OpenStudy (rsadhvika):

if x is the position function, say a projectile : -x^2+2x then, dx/dt is the first derivative, dx/dt = -2x+2

OpenStudy (rsadhvika):

fine wid that ?

OpenStudy (akashdeepdeb):

Yes. Kind of.

OpenStudy (rsadhvika):

d^2x/dt^2 is the second derivative, d^2x/dt^2 = d/dx(-2x+2) = -2 so the second derivative of poistion function(Acceleration) is -2

OpenStudy (akashdeepdeb):

Okay. Thanks but can you, by any chance, tell me what it means practically?

OpenStudy (rsadhvika):

we can take third derivative also for x :- d^3x/dt^2 = d/dt(-2) = 0

OpenStudy (rsadhvika):

okay, you comfortable wid notation part right ? d^2x/dt^2 is just a notation. dont do algebra with dx and dt...

OpenStudy (rsadhvika):

Next, lets try to relate dx/dt and d^x/dt^2 to velocity and acceleration

OpenStudy (rsadhvika):

say the position function x = -t^2 + 2t

OpenStudy (rsadhvika):

function x gives us exactly wat position the object is at, at any given time right ?

OpenStudy (rsadhvika):

this is the beginning

OpenStudy (anonymous):

What derivative means? Rate of change. So 'rate of change' of distance x is velocity v and 'rate of change' of velocity is acceleration...

OpenStudy (anonymous):

What does it MEAN?!

OpenStudy (amistre64):

the "derivative of a function" produces another function. since you can take the "derivative of a function" ... you can also take the (derivative of the "derivative of a function") ... all a derivative does is relate how the function is moving at a given point in the domain. When 2 function have the same successive derivatives all the way thru ... they move and act as one. if they move an act alike over all their domains, they are equivalent. If there is only a finite region in their domains that they act alike in, then their sameness is defined by the interval in which they converge.

OpenStudy (amistre64):

the rate of change in distance with respect to time is defined as velocity, velocity is a function of time the rate of change in velocity with respect to time is defined as acceleration, acceleration is a function of time the rate of change in acceleration with respect to time is defined as jerk, jerk is a function of time the rate of change of jerk with respect to time is to my knowledge not defined by a specific name; but it is still a function of time ...

OpenStudy (akashdeepdeb):

Thank you! I got quite a bit of that! But what does d mean in dx/dt And from that what would be d^2 signify?

OpenStudy (ybarrap):

d/dt of a function is just an operation on the function. You could just as well have wrote this as D(x) for 1st derivative or D(D(x)) to represent second derivative. If you think of this as just a function, then it might make more sense that D(D(x)) = \(\Large {d\over dt} {d\over dt}x={d^2x\over dt^2}\). The "d" and "dt" portions of this notation are derived from the ratios of "differences" between endpoints taken at smaller and smaller intervals: http://en.wikipedia.org/wiki/Derivative#Differentiation_and_the_derivative